# AgentOccam: A Simple Yet Strong Baseline for LLM-Based Web Agents

### v1

@@ Abstract
Autonomy via agents based on large language models (LLMs) that can carry out personalized yet standardized tasks presents a significant opportunity to drive human efficiency. There is an emerging need and interest in automating web tasks (e.g., booking a hotel for a given date within a budget). Being a practical use case itself, the web agent also serves as an important proof-of-concept example for various agent grounding scenarios, with its success promising advancements in many future applications. Meanwhile, much prior research focuses on handcrafting their web agent strategies (e.g., agent’s prompting templates, reflective workflow, role-play and multi-agent systems, search or sampling methods, etc.) and the corresponding in-context examples. However, these custom strategies often struggle with generalizability across all potential real-world applications. On the other hand, there has been limited study on the misalignment between a web agent’s observation and action representation, and the data on which the agent’s underlying LLM has been pre-trained. This discrepancy is especially notable when LLMs are primarily trained for language completion rather than tasks involving embodied navigation actions and symbolic web elements. In our study, we enhance an LLM-based web agent by simply refining its observation and action space, aligning these more closely with the LLM’s capabilities. This approach enables our base agent to significantly outperform previous methods on a wide variety of web tasks. Specifically, on WebArena, a benchmark featuring general-purpose web interaction tasks, our agent AgentOccam surpasses the previous state-of-the-art and concurrent work by 9.8 (+29.4%) and 5.9 (+15.8%) absolute points respectively, and boosts the success rate by 26.6 points (+161%) over similar plain web agents with its observation and action space alignment. We achieve this without using in-context examples, new agent roles, online feedback or search strategies. AgentOccam’s simple design highlights LLMs’ impressive zero-shot performance on web tasks, and underlines the critical role of carefully tuning observation and action spaces for LLM-based agents.
@@ 1 Introduction
AI agents leveraging large language models (LLMs) show great potential in automating repetitive and programmatic tasks and thereby alleviating human workloads (Gao et al., 2024; Xi et al., 2023; Yang et al., 2024). LLMs showcase remarkable capabilities in perception, reasoning and planning primarily due to their pre-training and post-learning. However, their effectiveness is significantly constrained when task-specific observation and action representations diverge from the parametric knowledge encoded during training of LLMs. For instance, in web-based tasks, these agents perform notably below human levels (Zhou et al., 2023b; Koh et al., 2024a).
@@
To improve web task performance by LLM-based agents, recent work focuses on designing better agent policies with either handcrafted prompting templates (Sodhi et al., 2024) or hard-coded auto-prompting strategies (Fu et al., 2024; Wang et al., 2024). While those pre-defined strategies can be effective for certain tasks, they struggle to generalize to diverse websites and varying skill requirements. Another emerging trend is to adopt sampling or search algorithms for a dynamic exploration of web navigation actions, which reduces dependence on pre-defined strategies but increases the cost of LLM inferences. (Koh et al., 2024b; Zhang et al., 2024; Pan et al., 2024).
@@
As shown in Figure 1, our method comprises of three components: i) We reduce non-essential actions to minimize the agent’s embodiment and trivial interaction needs; ii) We refine the observation by eliminating redundant and irrelevant web elements, and restructuring web content blocks for more succinct yet as informative representations; iii) We introduce two planning actions (branch and prune), which enables the agent to self-organize navigation workflow with a planning tree, and use the same structure to filter the previous traces for history replay. We implement these components by generic rules that applies to all types of markup-language-formatted web pages, without leveraging task-related information on the test benchmark.
By combining the three techniques mentioned above, our proposed agent AgentOccam performs substantially better on web tasks across websites in the WebArena environments (Zhou et al., 2023b)AgentOccam outperforms the previous state-of-the-art approach by 9.8 absolute points ($+29.4\%$) and surpasses concurrent work by 5.9 absolute points ($+15.8\%$). Notably, unlike most prior work, we do not use any in-context examples, additional online search or sampling, nor specialized prompting templates or agent roles to play well. In contrast, AgentOccam delivers such strong performance with an unexpectedly simple approach: letting the LLM issue actions within the processed and augmented observation and action spaces. Compared with a similar plain web agent without these proposed observation and action space changes, AgentOccam increases the success rate by 26.6 absolute points ($+161\%$).
@@
Essential Components Task-specific Strategies Additional Module In-context Examples Offline Data Online Search
@@
AWM (Wang et al., 2024) no yes yes yes 11footnotemark: 1 no
@@ 2 Related Work
Advances in large language and multi-modal foundation models have significantly boosted the development of autonomous agents to solve web tasks. Techniques translating LLMs to powerful decision-making agents (Yao et al., 2022b; Shinn et al., 2024) have led to progress in web agents, and have inspired many techniques that design inference time agent strategies. Many prior approaches improve the agent system by designing modular systems with specialized LLMs or roles, aiming to break down complex tasks (Sun et al., 2024; Prasad et al., 2024). Other works leverage LLMs to extract common patterns from examples or past experience (Zheng et al., 2023; Fu et al., 2024; Wang et al., 2024). However, this line of work often relies on pre-defined control hierarchy, prompt templates or examples to act accurately in the test environments. For example, SteP (Sodhi et al., 2024) utilizes a stack-based approach for dynamic multi-level control in the web tasks but relies on task-specific atomic policies with environment-related information hard-coded in prompt template. Another line of work focuses on improving web agents’ performance by leveraging more online examples from the environments. Many of them (Zhou et al., 2023a; Zhang et al., 2024; Putta et al., 2024) adapt Monte Carlo Tree Search (MCTS) methods, expanding intermediate states (tree nodes) in one task repeatedly by multiple trials over that task. Among them, WebPilot (Zhang et al., 2024) also adds a global optimization layer for high-level planning. Koh et al. (2024b) use a trained value function to guide search and to back-trace on the task execution tree. Auto Eval and Refine (Pan et al., 2024) trains a separate evaluator, and improves the task execution using reflective thinking (Shinn et al., 2024) on past trials in the same task. However, sampling or resetting multiple times in the same task, not only increases the inference cost significantly, but also limits its applicability when failed task is not revocable. As a comparison, we highlight the simplicity of our method and its difference with related agent approaches in Table 1.
@@
Fine-tuning language or multimodal models for web tasks is another effective approach to enhance decision-making capabilities on the web tasks (Yin et al., 2024; Hong et al., 2024; Lai et al., 2024; Putta et al., 2024). While fine-tuning promises more adaptivity and broader optimization space, the size of task-specific fine-tuned models are typically not comparable with the most powerful closed-source models. As for training models to follow natural language command on the computer or the web, there is also some early research before LLMs emerged, using semantic parsing (Artzi & Zettlemoyer, 2013), reinforcement learning (Branavan et al., 2009) and imitation learning (Liu et al., 2018; Humphreys et al., 2022). However, those fine-tuned agents, limited by the base model’s capacities or training data volume, often fail to match those constructed with LLMs regarding performance or/and generalizability, and is beyond the scope of this work.
@@
Web agent development has been supported by increasingly complex web simulators for training and evaluation. These range from basic platforms like MiniWoB (Shi et al., 2017) and its extension MiniWoB++ (Liu et al., 2018), to more sophisticated environments such as WebShop (Yao et al., 2022a), WebArena (Zhou et al., 2023b), and VisualWebArena (Koh et al., 2024a). These simulators progressively incorporate real-world complexities, from simple form-filling to tasks across multiple full-featured websites. In this work, we focus only on the text modality, and use WebArena to evaluate our method’s task success and generalizability as it contains different types of websites and task-intents in a single suite.
@@ 3 Problem Formulation
We formalize the web interaction process by a Partially Observable Markov Decision Process (POMDP, Littman (2009); Spaan (2012)): $\langle\mathcal{O},\mathcal{S},\mathcal{A},P,R,p_{0},\gamma\rangle$. In POMDPs, an observation $o\in\mathcal{O}$ consist of information that the agent receives from the web environment, e.g. HTMLs, as well as any instructions and prompts. In this work, we only consider the text modality. A state $s\in\mathcal{S}$ denotes the whole underlying (unobserved) state of the agent and the environment such that the state transition is Markovian. An action $a\in\mathcal{A}$ is either a command recognized by the web environment, or any other unrecognized token sequence that will lead to a stay in the current state. $P$ denotes a deterministic state transition function that records the change in the webpage state given the current state and agent action. $R$ is the reward function that decides the success or failure of the agent’s sequence of actions. In the WebArena environment used in our work, the reward is only assigned at the end of an agent-web interaction episode. $p_{0}$ denotes the initial state distribution which is uniform over 812 tasks in WebArena and discounting factor $\gamma$ is set to 1.
To solve POMDP, a common goal is to find a decision policy $\pi(a_{t}|h_{t})$ maximizing the expected cumulative reward, where $h_{t}$ denotes the observation history $\{o_{0},o_{1},...,o_{t}\}$. In LLM-based web agent design, that is translated to designing a policy $\pi(a_{t}|h_{t})$ with the help of one or more base LLM policy $\pi_{\textrm{LLM}}$ and a set of algorithmic modules. In this work, we work on a special class of policies that can be expressed as: $\pi(g(a_{t})|h_{t})=\pi_{\textrm{LLM}}(a_{t}|f(h_{t}))$, where $f$ and $g$ are rule-based functions that process the observation (including action instructions) and actions for the LLM policy. We name it the observation and action space alignment problem. Notice that under such problem setting, all of our changes apply only to the observations and the actions. We emphasize not all agent strategies in previous approaches can be represented in this way. For example: search-based algorithms require a control program on the top to select actions and trigger back-tracing; methods with evaluators, reflective thinking or memory modules also necessitate a managing center to alternate between the main LLM and these helper segments or other role-playing LLMs. In contrast, we aim to answer the following question in our work: Can we build a strong web agent with the base LLM policy $\pi_{\textrm{LLM}}$ by optimizing only the observation and action mapping $f$ and $g$?
@@ 4.1 Action Space Alignment
A web agent’s action space defines the valid commands it can use to interact with web environment. The WebArena simulator supports translating three categories of actions into mouse and keyboard operations: basic actions (e.g., click, type), tab operations (e.g., tab_focus for managing active tabs), and page operations (e.g., go_back for navigation). These actions are detailed in Appendix A, along with a comparison of our changes to the action space.
@@
Based on our observation of common failure modes in web agents, there are two key problems that need to be solved by editing the action space: i) removing irrelevant actions that LLMs struggle to understand and frequently misuse, and ii) improve the memorization and planning ability when the task execution requires navigating multiple potential paths to successfully execute. We propose that the first can be corrected simply by removing and combining actions. The second issue was often addressed in the previous work using handcrafted rules or strategies, making these approaches hard to generalize. In this work, we address the second problem by introducing actions that allow the LLM to autonomously generate plans and manage the task workflow. These proposed solutions are explained in details below and in Figure 2.
@@
Simplifying the Action Space. First, we eliminate actions that can be replicated using similar actions or replace multiple actions by one action with the same expressiveness (illustrated in Figure 2 step 1). Specifically, we remove the noop action, signifying “no operation”, as a distraction to the agent in most cases. Similarly, tab operations, which manage the focusing, opening, or closing of tabs are removed because they are only needed in a limited cases of multi-site tasks requiring two tabs. Furthermore, we limit page navigation actions like go_forward and goto, as their utility is greatly constrained by the agent’s poor memory of the relationship between a page’s URL and its content. By eliminating these less effective actions, our goal is to minimize distractions and boost the agent’s concentration on more meaningful operations. In addition, we introduce the note action, allowing the agent to record key observations for subsequent conclusions, and the stop action, enabling the agent to autonomously conclude the trajectory with answers. We also add a go_home command for multi-site tasks, enabling the agent to navigate directly to the homepage where all available sites are listed.
@@
Second, we eliminate actions that heavily require embodiment knowledge and simplify low-level actions into more abstract operations as shown in Figure 2 step 2. In particular, we reduce commands that LLM-based agents struggle with unless provided with detailed context-specific guidance, like hover or press (the latter is for pressing key combinations, often shortcuts). To properly use these actions requires LLMs to have embodied thinking of the current scenario, especially regarding the mouse position or keyboard operations, which it has not acquired during the training. Additionally, we remove the scroll action, opting instead to load the full page content as the web state. This change is in response to our observation that agents tend to engage in aimless and repetitive scrolling when an essential link is not visible at the top of the page, wasting steps without making progress. Furthermore, we streamline the agent’s interaction with drop-down menus; instead of selecting the menu and then an option, a single click command with the ID of desired option now suffices. The list of all actions in original and reduced action space are shown in Table 3, together with the frequency they are taken in different agents.
@@
Planning via Generation. Web tasks often requires solution that requires navigating multiple paths (e.g. extracting information from one page and submitting it to another page, like the task of creating a refund request on the contact us page for a broken product (task template 154), which requires parsing the order ID and refund amount from the order pages). We introduce two actions (branch and prune) to generate plans in a tree structure and save them for future observations. As Figure 2 step 3 shows, the LLM-generated plans starts with a root node being the objective of the task. The branch action will generate new subplans under the current node, decomposing high-level objectives into smaller, more manageable subgoals. Conversely, the prune action allows the agent to give up the current sub-plan (often after repetitive failed attempts) and seek for alternatives. Together with the branch and the prune actions, the LLM can edit the planning tree autonomously. Note that these two planning actions are of no difference from the native navigation actions in the web environment (e.g. click, type) and the LLM is free to choose when to take these actions to update the plan. The generated plan provides a context for future action generation and enhances the consistency of actions in one trajectory. This approach leverages the intrinsic planning ability in LLM itself. We argue that this will not compromise the agent’s generalization capability as this design relies minimally on prior knowledge.

### v2

@@ Abstract
Autonomy via agents based on large language models (LLMs) that can carry out personalized yet standardized tasks presents a significant opportunity to drive human efficiency. There is an emerging need and interest in automating web tasks (e.g., booking a hotel for a given date within a budget). Being a practical use case itself, the web agent also serves as an important proof-of-concept example for various agent grounding scenarios, with its success promising advancements in many future applications. Meanwhile, much prior research focuses on handcrafting their web agent strategies (e.g., agent’s prompting templates, reflective workflow, role-play and multi-agent systems, search or sampling methods, etc.) and the corresponding in-context examples. However, these custom strategies often struggle with generalizability across all potential real-world applications. On the other hand, there has been limited study on the misalignment between a web agent’s observation and action representation, and the data on which the agent’s underlying LLM has been pre-trained. This discrepancy is especially notable when LLMs are primarily trained for language completion rather than tasks involving embodied navigation actions and symbolic web elements. In our study, we enhance an LLM-based web agent by simply refining its observation and action space, aligning these more closely with the LLM’s capabilities. This approach enables our base agent to significantly outperform previous methods on a wide variety of web tasks. Specifically, on WebArena, a benchmark featuring general-purpose web interaction tasks, our agent AgentOccam surpasses the previous state-of-the-art and concurrent work by 9.8 ($+29.4\%$) and 5.9 ($+15.8\%$) absolute points respectively, and boosts the success rate by 26.6 points ($+161\%$) over similar plain web agents with its observation and action space alignment. Furthermore, on WebVoyager benchmark comprising tasks defined on real-world websites, AgentOccam exceeds the former best agent by 2.4 points ($+4.6\%$) on tasks with deterministic answers. We achieve this without using in-context examples, new agent roles, online feedback or search strategies. AgentOccam’s simple design highlights LLMs’ impressive zero-shot performance on web tasks, and underlines the critical role of carefully tuning observation and action spaces for LLM-based agents. [^1] Our code and data are available at https://github.com/amazon-science/AgentOccam.
@@ 1 Introduction
AI agents leveraging large language models (LLMs) show great potential in automating repetitive and programmatic tasks and thereby alleviating human workloads (Gao et al., 2024; Xi et al., 2023; Yang et al., 2024). LLMs showcase remarkable capabilities in perception, reasoning and planning primarily due to their pre-training and post-learning. However, their effectiveness is significantly constrained when task-specific observation and action representations diverge from the parametric knowledge encoded during their training/learning time. For instance, in web-based tasks, these agents perform notably below human levels (Zhou et al., 2023b; Koh et al., 2024a).
@@
To improve web task performance by LLM-based agents, recent work focuses on designing better agent policies with either handcrafted prompting templates (Sodhi et al., 2024) or hard-coded auto-prompting strategies (Fu et al., 2024; Wang et al., 2024). While those pre-defined strategies can be effective for certain tasks, they struggle to generalize to diverse websites and varying skill requirements. Another emerging trend is to adopt sampling or search algorithms for a dynamic exploration of web navigation actions, which reduces dependence on pre-defined strategies but increases the cost of LLM inferences (Koh et al., 2024b; Zhang et al., 2024; Pan et al., 2024).
@@
As shown in Figure 1, our method comprises of three components: i) We reduce non-essential actions to minimize the agent’s embodiment and trivial interaction needs; ii) We refine the observation by eliminating redundant and irrelevant web elements, and restructuring web content blocks for more succinct yet as informative representations; iii) We introduce two planning actions (branch and prune), which enables the agent to self-organize navigation workflow with a planning tree, and use the same structure to filter the previous traces for history replay. It’s noteworthy that those planning commands are in the same position as navigation prompts for the agent. We implement these components by generic rules that applies to all types of markup-language-formatted web pages, without leveraging task-related information on the test benchmark.
By combining the three techniques mentioned above, our proposed agent AgentOccam performs substantially better on web tasks across websites in the WebArena environments (Zhou et al., 2023b)AgentOccam outperforms the previous state-of-the-art approach by 9.8 absolute points ($+29.4\%$) and surpasses concurrent work by 5.9 absolute points ($+15.8\%$). Notably, unlike most prior work, we do not use any in-context examples, additional online search or sampling, nor specialized prompting templates or agent roles to play well. In contrast, AgentOccam delivers such strong performance with an unexpectedly simple approach: letting the LLM issue actions within the processed and augmented observation and action spaces. Compared with a similar plain web agent without these proposed observation and action space changes, AgentOccam increases the success rate by 26.6 absolute points ($+161\%$). Moreover, we prove AgentOccam’s web environment generalizability in the real-world web environment. In the WebVoyager definite-answer subset (He et al., 2024), which consists of real-world web tasks with deterministic answers, AgentOccam exceeds the previous best agent on this benchmark by 2.4 points ($+4.6\%$).
@@
Essential Components Task-specific Strategies In-context Examples Additional Module Offline Data [^2] The offline data refers to a labeled dataset to instill human knowledge into models. Online Search
@@
AWM (Wang et al., 2024) no yes yes yes  [^3] AWM supports two scenarios: in offline scenarios it directly leverage an offline dataset, and in online scenarios it relies on a domain-specific evaluator from Pan et al. (2024) which requires offline data to train. no
@@ 2 Related Work
Advances in large language and multi-modal foundation models have significantly boosted the development of autonomous agents to solve web tasks. Techniques translating LLMs to powerful decision-making agents (Yao et al., 2022b; Shinn et al., 2024) have advanced web agents, inspiring many techniques that design inference time agent strategies. Many prior approaches improve the agent system by designing auxiliary modules with specialized LLMs or roles, aiming to break down complex tasks (Sun et al., 2024; Prasad et al., 2024; Abuelsaad et al., 2024). Other work leverages LLMs to extract common patterns from examples or past experience (Zheng et al., 2023; Fu et al., 2024; Wang et al., 2024). However, this line of work often relies on pre-defined control hierarchy, prompt templates or examples to act accurately in the test environments. For example, SteP (Sodhi et al., 2024) utilizes a stack-based approach for dynamic multi-level control in the web tasks but relies on task-specific atomic policies with environment-related information hard-coded in prompt template. Another line of work focuses on improving web agents’ performance by leveraging more online examples from the environments. Many of them (Zhou et al., 2023a; Zhang et al., 2024; Putta et al., 2024) adapt Monte Carlo Tree Search (MCTS) methods, expanding intermediate states (tree nodes) in one task repeatedly by multiple trials over that task. Among them, WebPilot (Zhang et al., 2024) also adds a global optimization layer for high-level planning. Koh et al. (2024b) use a trained value function to guide search and to back-trace on the task execution tree. Auto Eval and Refine (Pan et al., 2024) trains a separate evaluator, and improves the task execution using reflective thinking (Shinn et al., 2024) on past trials in the same task. However, sampling or resetting multiple times in the same task, not only increases the inference cost significantly, but also limits its applicability when failed task is not revocable. As a comparison, we highlight the simplicity of our method and its difference with related agent approaches in Table 1.
@@
Fine-tuning language or multimodal models for web tasks is another effective approach to enhance decision-making capabilities on the web tasks (Yin et al., 2024; Hong et al., 2024; Lai et al., 2024; Putta et al., 2024). While fine-tuning promises more adaptivity and broader optimization space, the size of task-specific fine-tuned models are typically not comparable with the most powerful closed-source models. There is also some early research that trains models to follow natural language command on the web before LLMs emerged, using semantic parsing (Artzi & Zettlemoyer, 2013), reinforcement learning (Branavan et al., 2009) and imitation learning (Liu et al., 2018; Humphreys et al., 2022). However, those fine-tuned agents, limited by the base model’s capacities or training data volume, often fail to match those constructed with LLMs regarding performance or/and generalizability, and is beyond the scope of this work.
@@
Web agent development has been supported by increasingly complex web simulators for training and evaluation. These range from basic platforms like MiniWoB (Shi et al., 2017) and its extension MiniWoB++ (Liu et al., 2018), to more sophisticated environments such as WebShop (Yao et al., 2022a), WebArena (Zhou et al., 2023b), and VisualWebArena (Koh et al., 2024a). These simulators progressively incorporate real-world complexities, from simple form-filling to tasks across multiple full-featured websites. In this work, we focus only on the text modality, and use WebArena to evaluate our method’s task success and generalizability as it contains different types of websites and task-intents in a single suite. To further assess AgentOccam’s generalizabilty, we extend experiments to the real-world web environments, evluated with the tasks and golden answers proposed in WebVoyager (He et al., 2024).
@@ 3 Problem Formulation
We formalize the web interaction process by a Partially Observable Markov Decision Process (POMDP, Littman (2009); Spaan (2012)): $\langle\mathcal{O},\mathcal{S},\mathcal{A},P,R,p_{0},\gamma\rangle$. In POMDPs, an observation $o\in\mathcal{O}$ consists of information that the agent receives from the web environment, e.g. HTMLs, as well as any instructions and prompts. In this work, we only consider the text modality. A state $s\in\mathcal{S}$ denotes the whole underlying (unobserved) state of the agent and the environment such that the state transition is Markovian. An action $a\in\mathcal{A}$ is either a command recognized by the web environment, or any other unrecognized token sequence that will lead to a stay in the current state. $P$ denotes a deterministic state transition function that records the change in the webpage state given the current state and agent action. $R$ is the reward function that decides the success or failure of the agent’s sequence of actions. $p_{0}$ denotes the initial state distribution which is uniform over tasks tested. The discounting factor $\gamma$ is typically set to 1 when the reward is only assigned at the end of an agent-web interaction episode.
To solve POMDP, a common goal is to find a decision policy $\pi(a_{t}|h_{t})$ maximizing the expected cumulative reward, where $h_{t}$ denotes the observation history $\{o_{0},o_{1},...,o_{t}\}$. In LLM-based web agent design, that is translated to designing a policy $\pi(a_{t}|h_{t})$ with the help of one or more base LLM policies $\pi_{\textrm{LLM}}$ and a set of algorithmic modules. In this work, we work on a special class of policies that can be expressed as: $\pi(g(a_{t})|h_{t})=\pi_{\textrm{LLM}}(a_{t}|f(h_{t}))$, where $f$ and $g$ are rule-based functions that process the observation (including action instructions) and actions for the LLM policy. We name it the observation and action space alignment problem. Notice that under such a problem setting, all of our changes apply only to the observations and the actions. We emphasize not all agent strategies in previous approaches can be represented in this way. For example, search-based algorithms require a control program on the top to select actions and trigger back-tracing; methods with evaluators, reflective thinking or memory modules also necessitate a managing center to alternate between the main LLM and these helper segments or other role-playing LLMs. In contrast, we aim to answer the following question in our work: Can we build a strong web agent with the base LLM policy $\pi_{\textrm{LLM}}$ by optimizing only the observation and action mapping $f$ and $g$?
@@ 4.1 Action Space Alignment
A web agent’s action space defines the valid commands it can use to interact with the web environment. The WebArena simulator supports translating three categories of actions into mouse and keyboard operations: basic actions (e.g., click, type), tab operations (e.g., tab_focus for managing active tabs), and page operations (e.g., go_back for navigation). These actions are detailed in Appendix A, along with a comparison of our changes to the action space.
@@
Based on our observation of common failure modes in web agents, there exist two key challenges to be addressed by editing the action space: i) removing irrelevant actions that LLMs struggle to understand and frequently misuse, and ii) improving the memorization and planning ability when the task execution requires navigating multiple potential paths to succeed. We propose that the first can be corrected simply by removing and combining actions. The second issue was often addressed in the previous work using handcrafted rules or strategies, making these approaches hard to generalize. In this work, we tackle the second problem by introducing actions that allow the LLM to autonomously generate plans and manage the task workflow. These proposed solutions are explained in detail below and illustrated in Figure 2. The list of all actions in original and reduced action space is shown in Table 3, together with the frequency they are taken in different agents.
@@
Simplifying the Action Space. First, we eliminate actions that can be replicated using similar actions or replace multiple actions with one action with the same expressiveness (illustrated in Figure 2 step 1). Specifically, we remove the noop action, signifying “no operation”, as it is shown to be a distraction to the agent in most cases. Similarly, tab operations, which manage the focusing, opening, or closing of tabs are removed because they are only needed in limited cases of multi-site tasks requiring two tabs. Furthermore, we limit page navigation actions like go_forward and goto, as their utility is greatly constrained by the agent’s poor memory of the relationship between a page’s URL and its content. By eliminating these less effective actions, our goal is to minimize distractions and boost the agent’s concentration on more meaningful operations. In addition, we introduce the note action, allowing the agent to record key observations for subsequent conclusions, and the stop action, enabling the agent to autonomously conclude the trajectory with answers. We also add a go_home command for multi-site tasks, enabling the agent to navigate directly to the homepage where all available sites are listed.
@@
Second, we eliminate actions that heavily require embodiment knowledge and simplify low-level actions into more abstract operations as shown in Figure 2 step 2. In particular, we reduce commands that LLM-based agents struggle with unless provided with detailed context-specific guidance, like hover or press (the latter is for pressing key combinations, often shortcuts). To properly use these actions requires LLMs to have embodied thinking of the current scenario, especially regarding the mouse position or keyboard operations, which it has not acquired during their training. Additionally, we remove the scroll action, opting instead to load the full page content as the web state. This change is in response to our observation that agents tend to engage in aimless and repetitive scrolling when an essential link is not visible at the top of the page, wasting steps without making progress. Furthermore, we streamline the agent’s interaction with drop-down menus; instead of selecting the menu and then an option, a single click command with the ID of the desired option now suffices.
@@
Planning via Generation. Web tasks often require a solution that requires navigating multiple paths, e.g. extracting information from one page and submitting it to another page, like the task of creating a refund request on the Contact Us page for a broken product (task template 154), which requires parsing the order ID and refund amount from the order pages. We introduce two actions, branch and prune, to generate plans in a tree structure and save them for future observations. As Figure 2 step 3 shows, the LLM-generated plans start with a root node being the objective of the task. The branch action will generate new subplans under the current node, decomposing high-level objectives into smaller, more manageable subgoals. Conversely, the prune action allows the agent to give up the current sub-plan, often after repetitive failed attempts, and seek alternatives. Together with the branch and the prune actions, the LLM can edit the planning tree autonomously. Note that these two planning actions are of no difference from the native navigation actions in the web environment (e.g. click, type) and the LLM is free to choose when to take these actions to update the plan. The generated plan provides a context for future action generation and enhances the consistency of actions in one trajectory. This approach leverages the intrinsic planning ability of LLM itself. We argue that this will not compromise the agent’s generalization capability as this design relies minimally on task-specific prior knowledge.